40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 1 08:50:48 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial 
    Reply  Thu Jul 14 00:10:08 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial continued AS55_I.pngAS55_Q.pngREFL55_I.pngREFL55_Q.png
       Reply  Fri Jul 15 19:13:22 2022, Paco, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Fri Jul 15 22:36:04 2022, rana, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Mon Jul 18 16:39:07 2022, Paco, Summary, LSC, FPMI locking procedure using REFL55 and AS55 DARM_07_18_2022_FMPI.pdfCARM_07_18_2022_FPMI.pdfMICH_07_18_2022_FPMI.pdffpmi_darm_nb_2022_07.pdf
             Reply  Mon Jul 18 21:41:42 2022, Anchal, Summary, LSC, FPMI locking procedure using REFL55 and AS55 
             Reply  Tue Aug 9 19:54:31 2022, yuta, Summary, LSC, FPMI locking tonight 
                Reply  Thu Aug 18 14:49:35 2022, Yehonathan, Summary, LSC, FPMI Sensitivity Sensitivity_Plot_1344133503.pdf
                   Reply  Thu Aug 18 18:10:49 2022, Koji, Summary, LSC, FPMI Sensitivity 
Message ID: 17007     Entry time: Fri Jul 15 19:13:22 2022     In reply to: 17002     Reply to this: 17008   17012
Author: Paco 
Type: Summary 
Category: LSC 
Subject: FPMI with REFL/AS55 demod phase adjust 

[Yuta, Paco]

  • We first zero the offsets in ASDC, AS55, REFL55, POX11, and POY11 when PSL shutter is closed.
    • After this, we checked the offsets with only ITMX aligned. Some of RFPDs had ~2 counts of offsets, which indicate some RFAM of sidebands, but we decided not to tune Marconi frequencies since the offsets were small enough.
  • We went over the demod phases for AS55, REFL55, POX11, and POY11.
    • For POX11/POY11 first we just minimized the Q in each locked XARM/YARM individually. The newfound values were
      • C1:LSC-POX11_PHASE_R = 106.991
      • C1:LSC-POY11_PHASE_R = -12.820
    • Then we misaligned the XARM by getting rid of the MICH fringe in the ASDC port with ITMX yaw offset, and locked YARM using AS55_Q and REFL55_I and found the demod phase that minimized the AS55_I and REFL55_Q. The newfound values were
      • C1:LSC-AS55_PHASE_R = -65.9586
      • C1:LSC-REFL55_PHASE_R = -78.6254
    • Repeating the above, but now misaligning YARM with ITMY yaw offset, locking XARM with AS55_Q and REFL55_I, we found the demod phases that minimized AS55_1 and REFL55_Q. The newfound values were
      • C1:LSC-AS55_PHASE_R = -61.4361
      • C1:LSC-REFL55_PHASE_R = -71.0434
  • The above demod phases difference, Schnupp asymmetry between X and Y were measured. We repeated the measurement three times to derive the error.
    • Optimal demod phase difference between X arm and Y arm for both AS55 and REFL55 were measured to be -4.5 +/- 0.1 deg, which means that lx-ly = 3.39 +/- 0.05 cm (Marconi frequency: 11.066195 MHz).
  • We measured the gain difference between AS55_Q and POX11/POY11 = -0.5
  • We measured the gain difference between REFL55_I and POX11/POY11 = -2.5

After this, we locked DARM, CARM and MICH using POX11_I, POY11_I and AS55 error signals respectively, and actuating on ETMX, MC2, and BS with NO TRIGGERS (but FM triggers were on for boosts as usual). Under this condition, FM5 is used for lock acquisition, and FM1, FM2, FM3, FM6 are turned on with FM triggers. No FM4 was on. We also noticed:

  • CARM FM6 "BounceRoll" is slightly different than "YARM" FM6 "Bounce". The absent roll resonant gain actually makes it easier to control the CARM, we just had to use YARM filter for locking it.
  • When CARM is controlled, we often just kick the ETMX to bring it near resonance, since the frequency noise drops and we otherwise have to wait long.
ELOG V3.1.3-